OpenStudy (anonymous):

log7 1/343 help please

OpenStudy (anonymous):

\[\LARGE \log_7\left(\frac{1}{343}\right)=\] \[\LARGE \log_7\left(\frac{1}{7^3}\right)=\] \[\LARGE \log_7(7^{-3})=\] \[\LARGE -3\cdot \log_7(7)=\] \[\LARGE -3\]